if you want 5000 dollars in 3 years, what is your initial deposit if it is compounded continuously at 4.5%
Answer by rfer(16322)

A=Pe^rt
5000=P*2.7183^0.045*3
5000=P*1.1445
P=5000/1.1445
P=$4,368.72